# Learning parameter dependence for Fourier-based option pricing with tensor networks

### v2

@@ Acknowledgements.
We are grateful to Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the tensor cross interpolation code. We deeply appreciate Marc K. Ritter for his important remarks on the further improvement of our proposed method. R. S. is supported by the JSPS KAKENHI Grant No. 23KJ0295. R. S. thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing. R.S. is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka, as the inspiration for this study was derived from their collaboration. K. M. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant no. JPMXS0120319794, JSPS KAKENHI Grant no. JP22K11924, and JST COI-NEXT Program Grant No. JPMJPF2014.
@@ Appendix A How to set the tolerance of SVD
Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. The cutoff $\epsilon_{\text{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$. From Fig. 4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ while minimizing the computational complexity.

### v3

@@ Acknowledgements.
We are grateful to Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the tensor cross interpolation code. We deeply appreciate Marc K. Ritter and Hiroshi Shinaoka for their critical remarks on the further improvement of our proposed method. R. S. is supported by the JSPS KAKENHI Grant No. 23KJ0295. R. S. thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing. R.S. is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka, as the inspiration for this study was derived from their collaboration. K. M. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant no. JPMXS0120319794, JSPS KAKENHI Grant no. JP22K11924, and JST COI-NEXT Program Grant No. JPMJPF2014.
@@ Appendix A How to set the tolerance of SVD
Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{K}_{\mathrm{SVD}}$ to $10^{-10}$. The cutoff $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$. From Fig. 4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ while minimizing the computational complexity. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.